Mrs. Santos was 25 years old when her eldest daughter was born. Mrs. Santos' age now is 10 years more than 4 times her daughter's present age. What are their present ages?
I need an answer within an hour. I will be grateful if this will be answered. Thanks

Answer by stanbon(75887)

Mrs. Santos was 25 years old when her eldest daughter was born. Mrs. Santos' age now is 10 years more than 4 times her daughter's present age. What are their present ages?
-----------------
Present Age DATA:
Let daughter's age be "d"
Then mother's age is "4d+10"
----------------------------------------
Ages when daughter was born.
daughter was "0"
mother was (4d+10)-d = 3d+10
---------------------------------
Equation based on ages when daughter was born:
mother's age = 25
3d+10 = 25
3d = 15
d = 5 (daughter's age at present)
4d+10 = 20+10 = 30 (mother's age at present)
===========================
Cheers,
Stan H.